OpenStudy (anonymous):

Find the values of the trigonometric functions of t from the given information. sec t = 6, sin t < 0 im confused about doing this i can get cos t= 1/6 but then im stuck

OpenStudy (anonymous):

if you know cosine you can find sine

OpenStudy (anonymous):



OpenStudy (anonymous):

there is a picture of an angle where the cosine of the angle is \(\frac{1}{6}\) all that is missing is the other side of the triangle, which you can find via pyathagoras

OpenStudy (anonymous):

it is \[\sqrt{6^2-1^2}=\sqrt{35}\] and now you can find any trig ratio you like

OpenStudy (anonymous):

in particular \[\sin(t)=-\frac{\sqrt{35}}{6}\] the minus sign because you were told it is negative

OpenStudy (anonymous):

Thank you! i wasn't setting my triangle up right, i need to be more careful with that